# Bastani, Bastani, and McLaughlin

### v2

@@ 6 Conclusion
We have proposed a novel policy optimization strategy that aims to identify the full Pareto frontier of policies trading off expected performance from statistical significant performance improvement. Focusing on the specific setting where policy evaluation is performed using IPW, we have mathematically characterized the Pareto frontier. Our characterization reveals that focusing on statistical significance prioritizes elimination of bad treatment assignments over selecting the best treatment assignments. Intuitively, rather than try and identify optimal treatments, a decision-maker focuses on the problem of eliminating bad treatments, which can be simpler while yielding confident improvements. Based on our theoretical analysis, we devise a policy optimization algorithm using a plug-in approach. Practitioners first use their choice of supervised machine learning algorithm to train a counterfactual prediction model on the observational data, and then use its predictions to compute the Pareto frontier of optimal policies. Our simulations show that our plug-in approach can extract significant policy improvements, while other common policy optimization approaches cannot, especially when the underlying causal model is complex.
We leave a number of important directions for future work. First, while we have performed some preliminary analysis of how prediction errors affect estimation of the Pareto frontier, significantly more work is required to understand the impact. Our work also assumes the importance weights are known, whereas they may need to be estimated; understanding how IPW estimation error affects our procedure is another important direction. One key benefit of our convex formulation is that additional constraints (e.g., budget constraints) can be incorporated; it may be interesting to explore the impact of these constraints on the Pareto frontier. Finally, we have focused on IPW; exploring how to perform policy optimization in conjunction with other policy evaluation methodologies (e.g., randomized controlled trials, instrumental variables, etc.) deserves similar attention.

### v3

@@ 7 Conclusion
We have proposed a novel policy optimization strategy that aims to identify the full Pareto frontier of policies trading off expected performance with provably significant performance improvements. Focusing on the specific setting where policy evaluation is performed using IPW, we have mathematically characterized the Pareto frontier. Our characterization reveals that focusing on statistical significance prioritizes elimination of bad treatment assignments over selecting the best treatment assignments. Intuitively, rather than try and identify optimal treatments, a decision-maker focuses on the problem of eliminating bad treatments, which can be simpler while yielding confident improvements. Based on our theoretical analysis, we devise a policy optimization algorithm using a plug-in approach. Practitioners first use their choice of supervised machine learning algorithm to train a counterfactual prediction model on the observational data, and then use its predictions to compute the Pareto frontier of optimal policies. Our simulations show that our plug-in approach can extract significant policy improvements, while other common policy optimization approaches cannot, especially when the underlying causal model is complex.
We leave a number of important directions for future work. First, while we have performed some preliminary analysis of how prediction errors affect estimation of the Pareto frontier, significantly more work is required to understand the impact. Our work also assumes the importance weights are known, whereas they may need to be estimated; understanding how IPW estimation error affects our procedure is another important direction. One key benefit of our convex formulation is that additional constraints (e.g., budget constraints) can be incorporated; it may be interesting to explore the impact of these constraints on the characterization of the Pareto frontier. Finally, we have focused on IPW; exploring how to perform policy optimization in conjunction with other policy evaluation methodologies (e.g., randomized controlled trials, instrumental variables, etc.) deserves similar attention.